OpenStudy (anonymous):

angle OPE =(3x +170) angle EPN = (x^(2)) angle OPN = 180 . find x thanks!

OpenStudy (anonymous):

is this a quadrilateral, are you given a drawing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is a straight angle and a line coming out creating two more angles

OpenStudy (anonymous):

what letter is in the middle of the line?

OpenStudy (anonymous):

p sorry should have said that

OpenStudy (anonymous):

All we have to realize is that angleEPN +angleOPE = angle OPN

OpenStudy (anonymous):

Then you can set up an equation and then solve it.

OpenStudy (anonymous):

 not quite sure how to do that . should i set up the quadratic formula?

OpenStudy (anonymous):

i think that is what you end up with. simply write what the equation would be give angleEPN+angleOPE=angleOPN you are given expressions for all of these angles.

OpenStudy (anonymous):

ok just to be sure (3x +170 degrees) +x second power)degrees = 80 degrees?

OpenStudy (anonymous):

ok yu add them together and set them equal to 180

OpenStudy (anonymous):

was my above correct?( i meant to type 180 at the end)

OpenStudy (anonymous):

\[\left( 3x+170 \right)+x ^{2}=180\]

OpenStudy (anonymous):

ok thought so then would i use the quadratic equation to figure it out?

OpenStudy (anonymous):

\[x ^{2}+3x=10\]

OpenStudy (anonymous):

unless you want to set it to zero.

OpenStudy (anonymous):

thanks so much the helped lot!

OpenStudy (anonymous):

haha anytime.